OpenStudy (anonymous):

If f(x) = 5x2 – 3x + 4 and g(x) = 2x – 5, what is f(x) • g(x) ?

OpenStudy (anonymous):

Is simple as multiply both functions. f(x)*g(x)=(5x²-3x+4)(2x-5)

OpenStudy (anonymous):

Simplifiyng would be f(x)*g(x)=(10x³-25x²-6x²+15x+8x-20) Simplifying this f(x)*g(x)=(10x³-31x²+23x-5)

OpenStudy (anonymous):

you did something wrong it should either end with +20 or -20

OpenStudy (anonymous):

Yeah is -20 at the end i got mistaken.

OpenStudy (anonymous):

Thats ok